# Unveiling Downstream Performance Scaling of LLMs: A Clustering-Based Perspective

### v1

@@ 4.3 Extrapolation
We aim to identify clusters exhibiting robust scaling patterns for reliable performance extrapolation since some clusters have saturated or non-emergent performance on small models and are not expected to give reasonable predictions. We will show that performance prediction on these scalable clusters contributes to the prediction on the full evaluation set. We give the following definition to check whether a cluster is extrapolatable.
Definition: A cluster demonstrates scaling patterns if: (1) its expected accuracy increases monotonically with model size, and (2) the probability of correct responses converges to at least $P$ as computational resources approach infinity, where $P\leq 1$ is a predefined threshold accounting for practical limitations such as ambiguous questions and finite training coverage.
We filter clusters lacking scaling patterns by the following two rules:
@@
    Negligible accuracy growth with increased computational resources, manifested as minimal $a$ or $b$ values in Eq. 1;
@@
    Poor extrapolation reliability, indicated by excessive $c$ values in Eq. 1.
In practice, we set the parameter ranges a priori as $a>1$, $b>0.1$, and $10<c<$ 1.
The predictable subset comprises all samples from clusters that exhibit scaling patterns. The final performance prediction for a target model in the predictable subset is computed as the weighted average of the individual cluster predictions, with weights proportional to the cluster sizes.
@@ 4.4 Mapping from Predictable Subset to Target Evaluation Set
We extend our predictions from the predictable subset to the complete evaluation set through a principled mapping approach. Our method rests on the observation that extrapolatable and non-extrapolatable samples share question types but differ primarily in difficulty features, suggesting a preserved partial order of metrics across these subsets. We formalize this relationship through a mapping function $f:f(T^{\prime})\rightarrow T$ from predictable subset metrics $T^{\prime}$ to total evaluation set metrics $T$. This function exhibits key properties: 1. continuity and smoothness over $[0,1]$, 2. monotonic increase, and 3. passage through points $(0,0)$ and $(1,1)$. Empirical validation reveals that a quartic function optimally captures this relationship:
@@
To ensure reliable extrapolation, we calibrate the mapping curve using evaluation results of existing models as anchors. Our results show that the subset-to-full mapping generally maintains robustness across model architectures and training data, enabling the use of external models (e.g., Qwen2-72B [39]) as anchors for most tasks. We conduct corresponding experiments in Section 5.5. For data-sensitive tasks, models with similar training distributions provide more reliable anchors, indicating that data consistency takes precedence over architectural variation ensuring mapping accuracy. This calibration strategy enables accurate metric predictions for the complete evaluation set while maintaining computational efficiency.
Finally, combining Eq. 1 and Eq. 2, we get our final metric prediction $p=f\circ y(C_{0})$, where $C_{0}$ is the estimated computation of training the target LLM.
@@ 5.1 Experimental Setups
In our experimental setup, we train several smaller versions of the target model architecture for prediction. These models vary in size but share similar training procedures, with the training data scaled proportionally to their sizes.
Downstream evaluation sets. We adopt the following widely-used benchmarks as our target downstream tasks: For evaluation, we adopt the following widely-used benchmarks, shown in Table 1. Evaluation sets cover popular downstream tasks of the language model, including math, logic, coding, reading comprehension, professional knowledge, etc.
Table 1: Information of evaluation datasets used in the study.
Dataset GSM8K MATH BBH TriviaQA MBPP AGIEval DROP MMLU-pro
[7] [15] [32] [20] [3] [43] [9] [36]
Domain Math Math Reasoning Knowledge Coding Comprehensive Reading Comprehensive
#Questions 1,319 5,000 6,511 17,944 500 8,063 9,536 12,032
#Shots in Prompt 8 4 3 5 3 5 3 5
All models are evaluated in a few-shot in-context learning manner, where they need to generate final answer labels based on given demonstrations and test inputs. We aligned our evaluation setups with LLaMa3 [10].
Model training. To establish performance predictions for large language models, we conduct systematic experiments with a suite of smaller-scale models across different parameter counts and training data volumes, while controlling for other training configurations such as learning rate, batch size, and additional hyperparameters. All models are trained on a constant learning rate scheduler and data with the same distribution. We list model configurations in Table 2.
Table 2: Model architecture specifications across different sizes.
122M 238M 411M 652M 973M 1.9B 7B 12B 70B (Target)
Param. (M) 122 238 411 652 973 1,901 6,980 12,022 68,452
Tokens (B) 26 45 72 108 153 277 923 1,544 8,012
Layers 8 10 12 14 16 20 32 43 80
Model Dimension 1,024 1,280 1,536 1,792 2,048 2,560 4,096 4,608 8,192
Heads 8 10 12 14 16 20 32 36 64
KV Heads 8 10 12 14 16 20 8 12 8
@@ 5.2 Prediction Experiments
Baselines. We evaluate our proposed COD performance scaling for LLMs against existing approaches. The evaluation is conducted on multiple public benchmarks mentioned above, where we utilize a series of smaller models with identical data distribution and architecture but different configurations to estimate the downstream tasks performance of the target 70B large language model.
We compare against three existing prediction methods:
@@
    End-to-end performance-compute prediction: Extrapolate larger model metrics directly from smaller model evaluation set metrics using performance scaling laws.
@@
    Passrate-compute prediction: Estimate large model passrates from smaller model passrates [1, 19]. We repeat and evaluate 100 trials for each evaluation set to enhance the performance reliability on smaller models. For a fair comparison, we report the absolution prediction error on the passrate metric instead of greedy decoding accuracy.
@@
    Loss-intermediate performance prediction: First predict the final training loss of large language model, then estimate downstream task metrics based on the relationship between smaller model evaluation metrics and their corresponding losses [6].
We design two experimental groups to validate the benefits of clustering and the complete pipeline, respectively:
  1. 1.
    COD w/o. mapping: Performing difficulty-based clustering using K-Means, extrapolating within each cluster independently, and then aggregating metrics across clusters without requiring subset-to-full mappings.
  2. 2.
    COD complete: Complete multi-stage proposed approach consisting of clustering, predictable cluster filtering, subset extrapolation, and subset-to-full mapping.
The comparative results across different benchmarks and estimation approaches are presented in Table 3. We evaluate prediction accuracy with the absolute error between predicted and actual performance. We report the prediction error on each single evaluation set and list the mean and the max prediction error.
Table 3: Absolute prediction error on several evaluation sets. A prediction error less than 2 pp is considered an accurate estimate (marked in green), while an error greater than 5 pp is regarded as an invalid estimate (marked in red).
@@
Results. Predictions with an absolute error of less than 2 percentage points (pp) are considered accurate estimations, and when the predicted values fall within the training metric fluctuation range, they are marked in green; predictions with an absolute error greater than 5 indicate invalid estimations, which reduce the overall reliability of the prediction method and are marked in red. These results show our approach significantly outperforms existing methods in both mean and maximum prediction errors, maintaining mean prediction error within 2 pp, thus offering practical guidance for large model training. While existing methods demonstrate good performance on certain evaluation sets, they consistently exhibit substantial estimation errors on a minority of sets, undermining the credibility of their predictions.
Figure 4: Performance-compute relationship across difference prediction method
Through visualization of the performance-compute relationship, we illustrate the distinctive characteristics of different prediction methods, shown in Fig. 4. On the BBH evaluation set, while all three methods yield comparable estimates, end-to-end and loss-intermediate methods demonstrate inadequate fitting for small model evaluation points. In contrast, the cluster method reveals a more sophisticated and well-fitted multi-phase trajectory. For MATH and MMLU pro evaluation sets, the critical challenge involves determining whether large model metrics will experience accelerated growth with increased computing power or encounter performance plateaus. The loss-intermediate method exhibits an underestimation of the model capability ceiling, while the end-to-end method shows prediction errors exceeding 3 pp.
The clustering method’s effectiveness can be attributed to its comprehensive analysis of evaluation set difficulty distributions and scaling laws. It successfully predicts growth patterns in mathematical evaluation sets where most problems demonstrate expanded improvement potential, while accurately capturing the diminishing scaling properties in evaluation sets with score saturation as computational resources increase.
5.3 Comparison of Clustering Methods
We evaluate the impact of clustering methods on the estimation approach. Our goal is to control the average distance between samples and cluster centers within clusters, making difficulty features more similar within clusters. We also ensure that the minimum number of questions in any cluster is not less than 10, considering that too small clusters may lead to instability in metric values. We compared our proposed Improved-MeanShift algorithm with clustering methods including DBScan, MeanShift, and K-Means. Since standard K-Means lacks the ability to filter outliers and directly control intra-cluster distances, we made the following adjustments: (1) Search for the number of clusters such that the minimum cluster size is close to but not less than 10 samples; (2) Draw spheres around cluster centers with a given threshold radius, and treat samples not covered by any sphere as outliers. If a cluster drops to less than 10 samples, we treat its samples as outliers.
We use Intra-cluster Average Distance (IAD) and Outlier Rate (OR) as direct evaluation metrics. With similar OR, a smaller IAD indicates better clustering performance, as shown in Table 4. Additionally, we measure the benefits of different clustering methods on the prediction process by comparing the Extrapolation Errors(EE) of the predictable subset and Final prediction Errors (FE) after clustering, as shown in Table 5.
Table 4: Clustering performance on popular benchmarks.
@@ 5.3 Comparison of Clustering Methods
Table 5: Prediction errors across clustering algorithms.
@@
Table 4 shows that Improved-KMeans and Improved-MeanShift achieve better clustering performance, which is attributed to their incorporation of intra-cluster distance constraints during the clustering process. From Table 5, we can observe that the two methods with better clustering performance correspond to smaller extrapolation errors of estimable subsets and final metric prediction errors.
Although Improved-KMeans achieves optimal clustering performance, its downstream estimation performance on GSM8k is notably inferior compared to other evaluation sets. We believe this is because the K-Means algorithm requires pre-specifying a more explicit number of clusters, but the number of clusters in evaluation sets is difficult to know in advance. While the number of clusters obtained through our search method is effective for some evaluation sets, it lacks stability and ultimately leads to excessive prediction errors in a few evaluation sets. In contrast, our adopted Improved-MeanShift algorithm inherently does not require pre-specifying the number of clusters; instead, it automatically determines this based on our intra-cluster distance constraints. This results in a more stable clustering performance and yields the smallest maximum estimation error across evaluation sets. We present additional clustering experimental results on more evaluation sets in Section C.1, where the conclusions are consistent with existing evaluation sets.
5.4 Extrapolation Formula
To evaluate the effectiveness of different fitting formulas, we conducted an ablation study comparing various formulations of the accuracy-compute relationship. Our baseline formula incorporates random guess probability, exponential decay, and a constant offset term:
$f(C)=g+(1-g)*e^{-aC^{-b}-c}$ (3)
To understand the contribution of each component, we perform ablation experiments by removing or modifying different terms. 1) Without random guess component: $f_{1}(C)=e^{-aC^{-b}-c}$; 2) Without constant term $c$: $f_{2}(C)=g+(1-g)*e^{-aC^{-b}}$; 3) Direct power law relationship [19]: $f_{3}(C)=e^{-aC^{-b}}$.
The comparative results of these formulations are shown in Table 6. Results from three evaluation sets, BBH, Math, and MMLU-pro, are presented here, showing the Extrapolation Error of extrapolatable clusters (EE), the Task Ratio of predictable subset (TR), and the Final prediction Error (FE). These results show that the proposed formula $f$ consistently achieves the smallest extrapolation error and final prediction error, while the ratio of estimable subsets remains similar across different clustering methods.
Table 6: Ablation study results across different benchmarks.
@@ 5.4 Extrapolation Formula
Direct Power Law 8.90 49.06 8.88 3.81 81.46 3.35 4.30 95.15 4.27
w/o Random Guess 10.27 45.75 11.20 4.04 81.46 3.55 4.40 95.05 4.37
w/o Constant c 2.14 57.26 4.01 1.40 81.46 1.56 3.85 95.60 3.88
Ours 0.29 52.46 1.77 1.14 81.24 1.28 1.27 94.38 1.35
In the control group, $f_{1}$ performs poorly on tasks with finite answer sets, where small models achieve non-zero scores that $f_{1}$ cannot effectively fit. $f_{2}$ removes $c$, which determines the maximum value of the prediction curve, assuming that the evaluation set performance would reach perfect scores given sufficient computation and parameters. However, this assumption is unreasonable due to the limited training data distribution and ambiguous answers in evaluation set questions, leading to inaccurate predictions. Direct power-law fitting $f_{3}$ fails to model both the metric range constraint of 0 to 1 and the characteristic that metric improvement is typically more difficult near Random Guess capability and capability saturation compared to other regions.
We also observe that TR has little influence on prediction error, which also indicate the robustness of the proposed method. Some evaluation set with low TR due to the non-emergent subset, and our results show that this subset can largely be predicted with metrics of the extrapolatable clusters. We visualize the difficulty distribution of predictable subset and the full evaluation set in Appendix D.
5.5 Anchor Point in Interpolation Mapping
During the mapping phase from estimable subset metrics to full evaluation set metrics, we discovered that models with different training data and architectures exhibit similar mapping relationships. This allows us to leverage metrics from pre-trained models to refine the mapping relationship, thereby improving the accuracy of our final metric estimation.
We use both Qwen2-72B [39] and an in-house model $M$ with a Mixture-of-Experts(MoE) [22] structure trained on the same data distribution as anchor points in the mapping phase. We first obtain the interpolation curve using only small model metrics with points (0,0) and (1,1), then verify the compatibility of anchor points with the existing interpolation curve. When the score of the full set is 0 (1), the score of the subset must also be 0 (1). Results show that despite differences in computational requirements, model architectures, and training data, these models share similar mapping relationships.
This finding also indicates that estimable subset metrics are highly correlated with full-set metrics. Compared to loss-intermediate estimation, estimable subset metrics maintain predictability while reducing interference from other model parameters. Based on these observations, we incorporated the mapping relationships from pre-trained models into the interpolation process, thereby improving both the accuracy and confidence of our large model metric estimations.
We establish three experimental configurations of our method:
  • COD w/o. anchor: The complete estimation process is employed except for not using anchor point interpolation in the Mapping phase.
  • COD w. out-of-distribution (OOD) anchor: The complete proposed methodology incorporates both difficulty-based clustering and predictable subset identification. Using the 72B Qwen2 pretraining model as the anchor model [39].
  • COD w. in-domain(ID) anchor: Using an in-house MoE model with consistent training distribution but different model architecture as the anchor point.
We present the results in Table 7, demonstrating that incorporating both out-of-distribution models and in-distribution models as anchors consistently enhance prediction accuracy. These findings suggest a relatively stable correlation between the metrics of predictable subsets and the full dataset, indicating that the relationship between subset and full-set metrics remains consistent across models trained on different data and with varying architectures. This property enables us to leverage evaluation results from existing models to improve the accuracy of metric predictions for large models with new data and structures. Besides, the division of the evaluation set obtained through clustering is an intrinsic property of the evaluation set itself, independent of the model architecture and training data. Therefore, the predictable subset derived from clustering can also be extended to estimate the metrics of new models. Additionally, we conduct the ablation study on the interpolation method in Section C.2, and results indicate that quartic functions are suitable in our setting.
@@ 5.5 Anchor Point in Interpolation Mapping
Table 7: Influence of anchor point usage in the mapping stage.
@@
Mean Max GSM8k MATH BBH TriviaQA MBPP AGIEval DROP MMLU-pro

### v2

@@ 4.3 Extrapolation
To ensure reliable extrapolation, we identify clusters exhibiting robust scaling patterns, as some clusters may show saturated or non-emergent performance on smaller models, making them unsuitable for prediction.
A cluster is deemed extrapolatable if it meets two criteria: (1) its expected accuracy increases monotonically with model size, and (2) its performance converges to at least a predefined threshold $P$ (where $P\leq 1$ accounts for practical limits like ambiguous questions or finite training coverage).
We filter out non-extrapolatable clusters using two rules based on the parameters from Eq. 2:
@@
    Negligible accuracy growth, indicated by minimal $a$ or $b$ values.
@@
    Poor extrapolation reliability, indicated by an excessive $c$ value.
In practice, for extrapolatable clusters, we set $a>1$, $b>0.1$, and $10\leq c<$ 1.
The clusters satisfying these conditions form the predictable subset. The final performance prediction for a target model on this subset is the weighted average of the extrapolated predictions from these individual clusters, with weights proportional to cluster sizes.
@@ 4.4 Mapping from Predictable Subset to Target Evaluation Set
We map predictions from the predictable subset $T^{\prime}$ to the complete evaluation set $T$ using a function $f$. This approach is based on the observation that extrapolatable and non-extrapolatable samples, while differing in difficulty, often share question types, suggesting a preserved relative metric ordering. The mapping function $f:T^{\prime}\rightarrow T$ is continuous, smooth over $[0,1]$, monotonically increasing, and constrained to pass through $(0,0)$ and $(1,1)$. Empirical validation indicates a quartic function optimally captures this relationship:
@@
To ensure reliability, we calibrate $f$ using evaluation results from existing models as anchors. This subset-to-full mapping generally demonstrates robustness across diverse model architectures and training data, often permitting the use of external models (e.g., Qwen2-72B [40]) as anchors for many tasks (see Sec. 5.3.3 for experiments). For data-sensitive tasks, however, models with similar training distributions provide more reliable anchors, prioritizing data consistency for mapping accuracy. The final metric prediction for a target LLM with estimated training computation $C_{0}$ is then $p=f\circ y(C_{0})$, combining the cluster-wise extrapolation $y(C_{0})$ from Eq. 2 with the mapping $f$.
@@ 5.1 Experimental Setups
In our experimental setup, we train several smaller versions of the target model architecture for prediction. These models vary in size, but share similar training procedures, with the training data scaled proportionally to their sizes. We show the training configurations of the model in Appendix C.
For evaluation, we adopt the following widely used benchmarks, including GSM8K [7], MATH [16], BBH [33], TriviaQA [21], MBPP [3], AGIEval [44], DROP [9], MMLU-pro [36]. All models are evaluated in a few-shot in-context learning manner, and we aligned our evaluation setups with LLaMa3 [10]. We evaluate our proposed COD performance scaling for LLMs against existing approaches on multiple public benchmarks. Using a series of smaller models with identical data distribution and architecture but varied configurations, we estimate the downstream task performance of a pretrained LLM with 70B parameters.
@@ 5.2 Prediction Experiments
We compare COD against three representative prediction methods:
@@
    End-to-end performance-compute prediction: Directly extrapolates large model metrics from smaller model evaluation set metrics using performance scaling laws.
@@
    Passrate-compute prediction [1, 20]: Estimates large model passrates from smaller model passrates. We conduct 100 trials per evaluation set for smaller models to enhance reliability and report absolute prediction error on the passrate metric.
@@
    Loss-intermediate performance prediction [6]: First predicts the target LLM’s final training loss, then estimates downstream task metrics based on the relationship between smaller models’ evaluation metrics and their losses.
Table 1: Absolute prediction error (in percentage points) on evaluation sets. Errors < 2% are considered accurate (green), while errors > 5% are considered invalid (red). $\downarrow$ indicates lower is better.
@@
We also evaluate two variants of our COD approach to validate the benefits of its components:
  1. 1.
    COD (w/o mapping): Employs difficulty-based clustering, extrapolates within each cluster, and aggregates metrics, bypassing the subset-to-full mapping stage.
  2. 2.
    COD (Complete): Our full proposed multi-stage approach, including clustering, predictable cluster filtering, subset extrapolation, and subset-to-full mapping.
Comparative results are shown in Table 1. Prediction accuracy is measured by the absolute error between predicted and actual performance. We report mean and max prediction errors across all evaluation sets, as well as errors for individual sets. Our complete COD approach significantly outperforms existing methods in both mean (1.68%) and maximum (2.38%) prediction errors, offering reliable guidance for large model training. While baselines perform well on some sets, their substantial errors on others undermine overall credibility.
Figure 4: Performance-compute relationship for different prediction methods on BBH, MATH, and MMLU-pro evaluation sets.
Fig. 4 visualizes the performance-compute relationships. On BBH, while all methods give similar final estimates, end-to-end and loss-intermediate methods fit small model data poorly. COD reveals a more complex, better-fitted multi-phase trajectory. On MATH and MMLU-pro, where predicting accelerated growth versus plateaus is crucial, the loss-intermediate method underestimates model ceilings, and the end-to-end method exceeds 3% error. COD’s superior performance stems from its nuanced analysis of difficulty distributions and scaling laws, allowing it to predict growth in challenging sets (e.g., MATH) and capture diminishing returns in saturated sets.
The COD method is also capable of predicting LLMs’ performance after continual training phase on high-quality data. We conduct experiments and analysis on continual-trained LLMs and report prediction results in Appendix D.
5.3 Ablation Study
5.3.1 Comparison of Clustering Methods
We assess the impact of different clustering algorithms on prediction accuracy. The goal is to achieve tight intra-cluster difficulty similarity (low average distance to center) while maintaining cluster stability (min. 10 samples/cluster). We compare our Improved-MeanShift with DBScan, MeanShift, and K-Means. For K-Means, we adjust it to approximate our goals: (1) search for $k$ yielding min. cluster sizes $\approx 10$; (2) treat samples outside a radius (e.g., 2x average intra-cluster distance) from any cluster center as outliers, ensuring clusters don’t drop below 10 samples. We term this "Improved-KMeans" for this comparison. Clustering quality is measured by Intra-cluster Average Distance (IAD) and Outlier Rate (OR) (Table 2). Prediction benefits are measured by Extrapolation Error (EE) on the predictable subset and Final prediction Error (FE) (Table 3).
[…]
@@ 5.3.1 Comparison of Clustering Methods
[…]
@@
Table 2 shows Improved-KMeans and Improved-MeanShift yield better clustering (lower IAD) due to their intra-cluster distance constraints. Table 3 confirms these methods lead to lower EE and FE. Although Improved-KMeans has the best IAD, its GSM8k prediction is poor. This is likely because K-Means requires pre-specifying $k$, and our search for $k$ can be unstable, leading to large errors on some sets. In contrast, our Improved-MeanShift, which automatically determines $k$ based on distance constraints, offers more stable clustering and the lowest maximum prediction error. Additional results in Sec. E.1 support these conclusions.
5.3.2 Different Extrapolation Formulas
Table 4: Ablation study of extrapolation formulas. EE, TR, FE shown for BBH, MATH, MMLU-pro.
@@ 5.3.2 Different Extrapolation Formulas
Direct Power Law ($f_{3}$) 8.90 49.06 8.88 3.81 81.46 3.35 4.30 95.15 4.27
w/o Random Guess ($f_{1}$) 10.27 45.75 11.20 4.04 81.46 3.55 4.40 95.05 4.37
w/o Constant c ($f_{2}$) 2.14 57.26 4.01 1.40 81.46 1.56 3.85 95.60 3.88
Ours ($f$) 0.29 52.46 1.77 1.14 81.24 1.28 1.27 94.38 1.35
We ablate our proposed fitting formula $f(C)=g+(1-g)\cdot e^{-aC^{-b}-c}$ (Ours) by removing or modifying components: 1) $f_{1}(C)=e^{-aC^{-b}-c}$ (w/o random guess); 2) $f_{2}(C)=g+(1-g)\cdot e^{-aC^{-b}}$ (w/o constant $c$); 3) $f_{3}(C)=e^{-aC^{-b}}$ (Direct Power Law [20]). Table 4 shows Extrapolation Error (EE), Task Ratio of predictable subset (TR), and Final prediction Error (FE). Our proposed formula $f$ consistently achieves the lowest EE and FE. $f_{1}$ struggles with finite-answer tasks where small models have non-zero scores. $f_{2}$ inaccurately assumes perfect scores are attainable, ignoring data limitations and task ambiguities. The direct power law ($f_{3}$) fails to model the 0-1 metric range and the varying difficulty of improvement near random guess and saturation. TR’s minor influence on prediction error suggests robustness; even with low TR due to non-emergent tasks, their performance can be largely predicted via mapping (difficulty distributions visualized in Appendix F).
5.3.3 Different Anchor Points in Interpolation Mapping
We find that the mapping relationship from predictable subset metrics to full evaluation set metrics is similar across models with different training data and architectures. This allows leveraging pre-trained models as "anchors" to refine the mapping and improve final estimation accuracy. We use Qwen2-72B [40] (OOD anchor) and an in-house MoE model [23] with the same training data but different architecture (ID anchor). We first derive an interpolation curve using only small model metrics and fixed points $(0,0)$, $(1,1)$, then assess anchor compatibility. This shared mapping implies estimable subset metrics are highly correlated with full-set metrics and less prone to interference from other model parameters than loss-intermediate predictions.
We test three configurations:
  • COD (w/o anchor): Full COD pipeline, but no anchor points in the mapping phase;
  • COD (w. OOD anchor): COD with Qwen2-72B as an out-of-distribution anchor;
  • COD (w. ID anchor): COD with our in-house MoE model as an in-domain anchor.
@@ 5.3.3 Different Anchor Points in Interpolation Mapping
Table 5: Influence of anchor points in the mapping stage on prediction error (pp). Lower is better.
@@
[…]
@@
Table 5 shows that incorporating either OOD or ID anchors consistently enhances prediction accuracy. This suggests a stable correlation between predictable subset and full-set metrics across diverse models, enabling the use of existing model evaluations to improve predictions for new models. Furthermore, since our clustering identifies intrinsic properties of evaluation sets, the derived predictable subsets are applicable to new models. Additional ablations on the interpolation method (Sec. E.2) confirms quartic functions are suitable.